### v2

# Learning to Dissipate Energy in Oscillatory State-Space Models

Abstract
State-space models (SSMs) are a class of networks for sequence learning that benefit from fixed state size and linear complexity with respect to sequence length, contrasting the quadratic scaling of typical attention mechanisms. Inspired from observations in neuroscience, Linear Oscillatory State-Space models (LinOSS) are a recently proposed class of SSMs constructed from layers of discretized forced harmonic oscillators. Although these models perform competitively, leveraging fast parallel scans over diagonal recurrent matrices and achieving state-of-the-art performance on tasks with sequence length up to 50k, LinOSS models rely on rigid energy dissipation (“forgetting”) mechanisms that are inherently coupled to the time scale of state evolution. As forgetting is a crucial mechanism for long-range reasoning, we demonstrate the representational limitations of these models and introduce Damped Linear Oscillatory State-Space models (D-LinOSS), a more general class of oscillatory SSMs that learn to dissipate latent state energy on arbitrary time scales. We analyze the spectral distribution of the model’s recurrent matrices and prove that the SSM layers exhibit stable dynamics under a simple, flexible parameterization. Without introducing additional complexity, D-LinOSS consistently outperforms previous LinOSS methods on long-range learning tasks, achieves faster convergence, and reduces the hyperparameter search space by 50%.
1 Introduction
State-space models (SSMs) (Gu et al., 2021; Smith et al., 2023; Gu & Dao, 2023; Hasani et al., 2022; Rusch & Rus, 2025) have emerged as a powerful deep learning architecture for sequence modeling, demonstrating strong performances across various domains, including natural language processing (Gu & Dao, 2023), audio generation (Goel et al., 2022), reinforcement learning (Lu et al., 2024), and scientific and engineering applications (Hu et al., 2024).
Despite the abundance of neural network architectures for sequence modeling, SSMs have gained significant attention due to their fundamental advantages over both Recurrent Neural Networks (RNNs) and Transformer architectures based on self-attention mechanisms (Vaswani, 2017). Built upon layers of sequence-to-sequence transformations defined by linear dynamical systems, SSMs integrate principles from control theory with modern deep learning techniques, making them highly effective across multiple modalities. While recent SSM architectures are often formulated as linear RNNs (Orvieto et al., 2023), they introduce notable improvements over their predecessors, offering enhanced speed, accuracy, and the ability to capture long-range dependencies more effectively.
In this work, we focus on the regime of linear, time-invariant (LTI) SSMs and extend the recently introduced Linear Oscillatory State-Space model (LinOSS) (Rusch & Rus, 2025). LinOSS formulates a continuous-time system of second-order ordinary differential equations (ODEs) that represent forced harmonic oscillators. These dynamics are then discretized into two conditionally stable state-space variants, each derived from a different ODE integration method, and the resulting models are computed efficiently by leveraging associative parallel scans. The structure of the underlying oscillatory dynamics allows LinOSS to learn long-range interactions with minimal constraints on the SSM state matrix. However, previous LinOSS models inherently couple the frequency and damping behaviors of state-space layers, effectively collapsing latent state energy dissipation to a single scale and limiting the model’s expressivity. To overcome this, we introduce a flexible and controllable dissipation mechanism and derive the Damped Linear Oscillatory State-Space Model (D-LinOSS), enhancing the LinOSS architecture by incorporating learnable damping independent of time scale.
Our approach constructs a deep state space model capable of capturing a wide range of temporal relationships by expanding the expressivity of individual SSM layers. Unlike previous versions of LinOSS that were constrained to a limited subset of oscillatory systems, our method allows each layer to independently learn a wider range of stable oscillatory dynamics, collectively leading to a more powerful sequence model. Our full contributions are:
  • We conduct a rigorous spectral analysis of the proposed D-LinOSS model, highlighting the representational improvements enabled by learnable damping.
  • We validate the theoretical expressivity improvements through a synthetic experiment of learning exponential decay.
  • We derive a stable parameterization of D-LinOSS and introduce an initialization procedure to generate arbitrary eigenvalue distributions in the recurrent matrix. We perform ablations comparing different initialization techniques.
  • We provide extensive empirical evaluation, showing that D-LinOSS on average outperforms state-of-the-art models across eight different challenging real-world sequential datasets.
  • We showcase the additional practical benefits of D-LinOSS, such as faster convergence time and a 50% reduction in model hyperparameter space compared to previous LinOSS models.
  • To support reproducibility and further research, we release our code and experiments at github.com/jaredbmit/damped-linoss.
Figure 1: Previous LinOSS models directly couple the frequency and magnitude of discretized system eigenvalues, reducing latent state energy dissipation to a single scale when normalizing time by frequency. Through a simple, complexity-free parametric extension, D-LinOSS can learn system damping on any scale, independent of frequency, expanding the range of expressible internal dynamics. The specific damping behavior depicted in the right diagram is selected arbitrarily.
2 Background
2.1 Continuous-time formulation
D-LinOSS layers are constructed from a system of damped, forced harmonic oscillators:
$\mathbf{x}^{\prime\prime}(t)$ $=-\mathbf{A}\mathbf{x}(t)-\mathbf{G}\mathbf{x}^{\prime}(t)+\mathbf{B}\mathbf{u}(t),$ (1)
$\mathbf{y}(t)$ $=\mathbf{C}\mathbf{x}(t)+\mathbf{D}\mathbf{u}(t)$
The continuous-time parameters $\mathbf{A}$ and $\mathbf{G}$ are restricted to diagonal matrices with non-negative entries, meaning (1) is an uncoupled second-order system. The feed-forward operation $\mathbf{D}\mathbf{u}(t)$ will be omitted for the rest of the paper for concision.
D-LinOSS layers provide an expressive, stable, and efficient recurrent primitive for modeling intermediate sequence-to-sequence transformations $\mathbf{u}\mapsto\mathbf{y}$ in $\mathbb{R}^{m}$ through learning parameters $\mathbf{A}$, $\mathbf{G}$, $\mathbf{B}$, and $\mathbf{C}$. $\mathbf{A}$ controls the natural frequency of the system’s oscillation and $\mathbf{G}$ defines the damping, i.e., the energy dissipation of the latent state. The underlying dynamical system of previous LinOSS models is (1) subject to $\mathbf{G}=\mathbf{0}$; thus, D-LinOSS is constructed from a more general oscillatory dynamical system with learnable damping. The additional $m$ learnable parameters from $\mathbf{G}$ are a negligible contribution to model size and have no impact on speed.
2.2 Discretization
The D-LinOSS discrete-time state-space layer is derived by adopting an integration scheme to approximately solve System (1) as an initial-value problem (IVP) subject to $\mathbf{x}(0)=\mathbf{x}^{\prime}(0)=\mathbf{0}$. This discretization technique is effectively a specification for mapping the underlying continuous-time parameters $\mathbf{A},\mathbf{G},$ and $\mathbf{B}$ to discrete-time counterparts $\mathbf{M}$ and $\mathbf{F}$ in the following systems.
$\begin{bmatrix}\mathbf{z}^{\prime}(t)\\ \mathbf{x}^{\prime}(t)\end{bmatrix}$ $=\begin{bmatrix}-\mathbf{G}&-\mathbf{A}\\ \mathbf{I}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{z}(t)\\ \mathbf{x}(t)\end{bmatrix}+\begin{bmatrix}\mathbf{B}\\ \mathbf{0}\end{bmatrix}\mathbf{u}(t)$ (2) $\longrightarrow$ $\begin{bmatrix}\mathbf{z}_{k+1}\\ \mathbf{x}_{k+1}\end{bmatrix}$ $=\mathbf{M}\begin{bmatrix}\mathbf{z}_{k}\\ \mathbf{x}_{k}\end{bmatrix}+\mathbf{F}\mathbf{u}_{k+1}$ (3)
Discretization also introduces learnable time-step parameters $\Delta t\in\mathbb{R}^{m}$ that govern the integration interval for the ODE solution.
Unlike standard first-order SSMs, oscillatory SSMs explicitly model the acceleration and velocity of the system state, resulting in smoother outputs due to the twice-integrated dynamical structure. As a result, although most SSMs discretize the continuous-time dynamics using zero-order hold or the bilinear method, the second-order structure of D-LinOSS necessitates the use of special discretization schemes to maintain conditional system stability without over-constraining the matrices $\mathbf{A}$ and $\mathbf{G}$.
Specifically, Rusch & Rus (2025) investigate the use of implicit integration (IM) and symplectic integration, also referred to as implicit-explicit integration (IMEX), as discretization methods for their proposed continuous-time system of undamped harmonic oscillators. Each integrator endows the resulting SSM with different energy dissipation properties; IM integration produces a dissipation term coupled to the eigenvalue phase and IMEX integration completely preserves energy across time. The selection of discretization technique is thus a binary hyperparameter in the original LinOSS model used to modulate the amount of “forgetting," giving rise to two SSMs (LinOSS-IM and LinOSS-IMEX) exhibiting different dynamical behaviors.
We extend the use of IMEX integration to the D-LinOSS framework, as learnable damping allows for full dynamical control regardless of which discretization method is used. This flexibility in parameterization removes the need to treat discretization schemes as a hyperparameter, reducing the search space by 50%.
Applying IMEX integration to System (1) yields:
$\mathbf{z}_{k+1}$ $=\mathbf{z}_{k}+\Delta t\big(-\mathbf{A}\mathbf{x}_{k}-\mathbf{G}\mathbf{z}_{k+1}+\mathbf{B}\mathbf{u}_{k+1}\big),$ (4)
$\mathbf{x}_{k+1}$ $=\mathbf{x}_{k}+\Delta t\mathbf{z}_{k+1}$
or in matrix form,
$\begin{bmatrix}\mathbf{I}+\Delta t\mathbf{G}&\mathbf{0}\\ -\Delta t\mathbf{I}&\mathbf{I}\end{bmatrix}\begin{bmatrix}\mathbf{z}_{k+1}\\ \mathbf{x}_{k+1}\end{bmatrix}=\begin{bmatrix}\mathbf{I}&-\Delta t\mathbf{A}\\ \mathbf{0}&\mathbf{I}\end{bmatrix}\begin{bmatrix}\mathbf{z}_{k}\\ \mathbf{x}_{k}\end{bmatrix}+\begin{bmatrix}\Delta t\mathbf{B}\\ \mathbf{0}\end{bmatrix}\mathbf{u}_{k}$ (5)
Inverting the left hand side block matrix, we arrive at the final discrete-time SSM in the form of (3).
$\mathbf{M}:=\begin{bmatrix}\mathbf{S}^{-1}&-\Delta t\mathbf{S}^{-1}\mathbf{A}\\ \Delta t\mathbf{S}^{-1}&\mathbf{I}-\Delta t^{2}\mathbf{S}^{-1}\mathbf{A}\end{bmatrix},\quad\mathbf{F}:=\begin{bmatrix}\Delta t\mathbf{S}^{-1}\mathbf{B}\\ \Delta t^{2}\mathbf{S}^{-1}\mathbf{B}\end{bmatrix}$ (6)
Here, the Schur complement is the diagonal matrix $\mathbf{S}=\mathbf{I}+\Delta t\mathbf{G}$ and $\mathbf{M}$ is a block matrix composed of diagonal sub-matrices.
2.3 Associative parallel scans
Many modern SSM architectures (Smith et al., 2023) leverage associative parallel scans (Kogge & Stone, 1973; Blelloch, 1990) to efficiently compute recurrent operations across long sequences. By exploiting the associativity of the recurrence operator, naively $O(N)$ sequential operations can be parallelized and computed in $O(\log N)$ time. For SSMs, parallel scans enable sub-linear complexity of the recurrence computation with respect to sequence length, acting as a key building block for scaling SSMs to long contexts.
3 Theoretical properties
Spectral analysis provides a lens to examine the stability and dynamical behavior of SSMs. In the absence of bounding nonlinearities like $\tanh$, the eigenvalues of the recurrent matrix $\mathbf{M}$ fully govern how latent states evolve across time. In particular, eigenvalues with near unit norm retain energy across long time horizons, while those closer to zero rapidly dissipate energy.
In the previous LinOSS-IM and LinOSS-IMEX models, the internal system spectra are rigidly defined by the selection of discretization technique, coupling frequency and damping. As shown in Figure 1, this effectively reduces latent state energy dissipation to a single scale, limiting the range of expressible dynamics. For D-LinOSS, the spectrum of $\mathbf{M}$ instead arises from damped harmonic oscillators, introducing a new tunable mechanism that decouples damping from frequency. Unlike the preceding models, D-LinOSS layers can represent all stable second-order systems, yielding a broader range of expressible dynamics and thus a more powerful sequence model. Figure 1 depicts this, where the scale of energy dissipation can be arbitrarily selected regardless of oscillation frequency.
These notions are formalized in this section, where we characterize the eigenvalues of D-LinOSS, derive stability conditions, and compare the resulting spectral range to that of previous LinOSS models. In particular, we show that the set of reachable, stable eigenvalue configurations in D-LinOSS is the full complex unit disk, where that of LinOSS has zero measure in $\mathbb{C}$.
3.1 Spectral analysis and stability
Proposition 3.1.
The eigenvalues of the D-LinOSS recurrent matrix $\mathbf{M}\in\mathbb{R}^{2m\times 2m}$ are
$\lambda_{i_{1,2}}=\frac{1+\frac{\Delta t_{i}}{2}\mathbf{G}_{i}-\frac{\Delta t_{i}^{2}}{2}\mathbf{A}_{i}}{1+\Delta t_{i}\mathbf{G}_{i}}\pm\frac{\frac{\Delta t_{i}}{2}\sqrt{(\mathbf{G}_{i}-\Delta t_{i}\mathbf{A}_{i})^{2}-4\mathbf{A}_{i}}}{1+\Delta t_{i}\mathbf{G}_{i}},$ (7)
where pairs of eigenvalues are denoted as $\lambda_{i_{1,2}}$ and i=1,2,,mi=1,2,...,m.
Proof.
The derivation is provided in Appendix A.1. Because the $m$ second-order systems are decoupled, it is sufficient to subsequently analyze the spectral properties and stability conditions for a single system with index $i\in{1,\dots,m}$. ∎
Proposition 3.1 shows that eigenvalues of D-LinOSS are tuned through choices of $\mathbf{A}$, $\mathbf{G}$, and $\Delta t$. We now detail a sufficient condition for system stability.
Proposition 3.2.
Assume $\mathbf{A}_{i}$, $\mathbf{G}_{i}$ are non-negative, and $\Delta t_{i}\in(0,1]$. If the following is satisfied:
$(\mathbf{G}_{i}-\Delta t_{i}\mathbf{A}_{i})^{2}\leq 4\mathbf{A}_{i},$ (8)
then $\lambda_{i_{1,2}}$ come in complex conjugate pairs $\lambda_{i}$, $\lambda_{i}^{*}$ with the following magnitude:
$|\lambda_{i}|=\frac{1}{\sqrt{1+\Delta t_{i}\mathbf{G}_{i}}}\leq 1,$ (9)
i.e., the eigenvalues are unit-bounded. Define $\mathcal{S}_{i}$ to be the set of all $(\mathbf{A}_{i},\mathbf{G}_{i})$ that satisfy the above condition. For notational convenience, we order the eigenvalues such that $\text{\emph{Im}}(\lambda_{i})\geq 0$, $\text{\emph{Im}}(\lambda_{i}^{*})\leq 0$.
Proof.
The proof is detailed in Appendix A.2. Condition (8) is simply the non-positivity of the discriminant in the eigenvalue expression of Proposition 3.1, which is shown to be sufficient for the unit-boundedness of $|\lambda_{i}|$. ∎
We now demonstrate that the spectral image of $\mathcal{S}_{i}$ is the full unit disk, meaning D-LinOSS is capable of representing every stable, damped, uncoupled second-order system.
Proposition 3.3.
The mapping $\Phi:\mathcal{S}_{i}\rightarrow\mathbb{C}_{|z|\leq 1}$ defined by $(\mathbf{A}_{i},\mathbf{G}_{i})\mapsto(\lambda_{i},\lambda_{i}^{*})$ is bijective.
Proof.
In Appendix A.3, a well-defined inverse mapping $\Phi^{-1}:\mathbb{C}_{|z|\leq 1}\rightarrow\mathcal{S}_{i},\ (\lambda_{i},\lambda_{i}^{*})\mapsto(\mathbf{A}_{i},\mathbf{G}_{i})$ is constructed. This inverse map has practical utility in matrix initialization, enabling the selection of arbitrary distributions of stable eigenvalues. ∎
In contrast to the full expressive range of D-LinOSS layers, LinOSS-IM and LinOSS-IMEX layers are limited in reachable eigenvalues. We recall the respective expressions from Rusch & Rus (2025):
$\lambda^{\text{IM}}_{i_{1,2}}=\frac{1}{1+\Delta t_{i}^{2}\mathbf{A}_{i}}\pm j\frac{\Delta t_{i}\sqrt{\mathbf{A}_{i}}}{1+\Delta t_{i}^{2}\mathbf{A}_{i}},\quad\lambda^{\text{IMEX}}_{i_{1,2}}=\frac{1}{2}(2-\Delta t_{i}^{2}\mathbf{A}_{i})\pm\frac{j}{2}\sqrt{\Delta t_{i}^{2}\mathbf{A}_{i}(4-\Delta t_{i}^{2}\mathbf{A}_{i})},$ (10)
It can be seen that both forms impose a rigid relationship between frequency and damping, constraining the reachable spectra. The following proposition formalizes this by showing that the set of eigenvalues reachable under these parameterizations occupies zero area within the unit disk. Appendix A.4 shows the respective expressions graphed on $\mathbb{C}_{|z|\leq 1}$ for a visual interpretation.
Proposition 3.4.
For both LinOSS-IM and LinOSS-IMEX, the set of eigenvalues constructed from $\mathbf{A}_{i}\in\mathbb{R}_{\geq 0}$ and $\Delta t_{i}\in(0,1]$ is of measure zero in $\mathbb{C}$.
Proof.
Detailed in Appendix A.4, the change of variables $\gamma_{i}=\Delta t_{i}\sqrt{\mathbf{A}_{i}}$ renders both eigenvalue expressions one-dimensional curves, which have zero measure in $\mathbb{C}$. ∎
The incorporation of explicitly learnable damping enables D-LinOSS to model a wider range of stable dynamical systems, increasing the representational capacity of the overall deep sequence model. The empirical performance benefits are discussed in the following sections.
3.2 Motivation
A natural question is whether or not a larger set of reachable eigenvalues is empirically useful. Notably, LinOSS is provably universal (Rusch & Rus, 2025; Lanthaler et al., 2024), meaning it can approximate any causal and continuous operator between input and output signals to arbitrary accuracy (see Appendix A.5 for a formal definition). This property trivially extends to D-LinOSS, as setting $\mathbf{G}=\mathbf{0}$ recovers LinOSS-IMEX. However, while universality characterizes theoretical capacity, it is not necessarily indicative of how well a model learns in practice. To motivate the empirical benefits of a broader spectral range, we investigate the following synthetic experiment.
Table 1: Learning exponential decay.
Model RMSE $\times 10^{-3}$
LinOSS-IMEX 24.5 $\pm$ 2.3 30.6 $\times$
LinOSS-IM 8.0 $\pm$ 1.7 10.0 $\times$
D-LinOSS 0.8 $\pm$ 0.1 1.0 $\times$
We simulate a dynamical system with a single discrete eigenvalue $\lambda=0.8$, corresponding to an exponentially decaying response. No input or output transformations are applied. Random sequences of scalar values are passed through this system, and models are trained to predict the resulting output. A small hyperparameter grid was searched for each model, three random seeds were trained per configuration, and models are compared in terms of the best average test root mean squared error (RMSE) across seeds. (Appendix B.1 provides more detail on the experimental setup.)
D-LinOSS achieves test RMSE approximately $10\times$ lower than LinOSS-IM and $30\times$ lower than LinOSS-IMEX. Reflecting on the eigenvalue ranges depicted in Appendix A.4, the target eigenvalue $\lambda=0.8$ lies outside the reachable spectra of both baseline models. This gap in performance suggests that although the previous LinOSS models are universal, a limited spectral range can impair the models’ ability to represent certain temporal relationships. By directly learning system damping $\mathbf{G}$, D-LinOSS moves beyond rigidly defined energy dissipation behavior and is capable of accurately capturing a wider range of dynamics.
3.3 Parameterization
During training, the learnable parameters $\mathbf{A},\mathbf{G}$, and $\Delta t$ must satisfy the constraints detailed in Proposition 3.2 to ensure system stability. Unconstrained (denoted by bar) continuous-time SSM matrices and integration time steps are reparameterized through the following activations to enforce these conditions.
$\Delta t=\sigma(\bar{\Delta t}),\quad\mathbf{G}=\text{ReLU}(\bar{\mathbf{G}}),\quad\mathbf{A}=\text{Clamp}(\bar{\mathbf{A}},L(\mathbf{G},\Delta t),U(\mathbf{G},\Delta t))$
Here, $\text{Clamp}(\cdot)$ bounds $\bar{\mathbf{A}}$ between lower bound $L$ and upper bound $U$, which are expressions derived from the quadratic inequality of the stability criterion $\mathcal{S}$.
3.4 Initialization
Performance of SSMs can be heavily impacted by the initialized distribution of eigenvalues (Orvieto et al., 2023). Many approaches, e.g., Gu et al. (2021); Smith et al. (2023), leverage structured initialization schemes such as the HiPPO framework (Gu et al. (2020)) to enhance long-range learning capability, but Orvieto et al. (2023) indicates that simpler initialization techniques, such as uniform ring initialization, are capable of recovering equal performance.
Given the parameterization in Section 3.3 and the bijective mapping between the parameter space $(\mathbf{A}_{i},\mathbf{G}_{i})$ and the eigenspace $(\lambda_{i},\lambda_{i}^{*})$ established in Proposition 3.3, we investigate initialization strategies for SSM layers based on uniform sampling in each of these spaces. Appendix B.5 details a comparison between two approaches: (i) uniformly sampling parameters $(\mathbf{A}_{i},\mathbf{G}_{i})$ across different ranges $A_{\text{min}},A_{\text{max}}]$ and $G_{\text{min}},G_{\text{max}}]$, and (ii) uniformly sampling eigenvalues within a complex ring across different radial bounds $r_{\text{min}},r_{\text{max}}]$ and angular bounds $\theta_{\text{min}},\theta_{\text{max}}]$. In the latter case, parameters are then obtained via the inverse mapping $(\mathbf{A}_{i},\mathbf{G}_{i})=\Phi^{-1}(\lambda_{i})$ for all $i\in\{1,\dots,m\}$. Our study indicates that sampling eigenvalues within the radial band $[0.9,1.0]$ and the full angular range $0,2\pi)$ yields strong performance. This technique is used for all subsequent D-LinOSS results.
4 Results
Following the experimental design from Rusch & Rus (2025); Walker et al. (2024); Zhou et al. (2021), we evaluate the empirical performance of D-LinOSS on a suite of real-world learning tasks that span disciplines across biology, medicine, chemistry, photonics, and climate. As the linear complexity and fixed state size of SSMs emphasize their utility for learning long-range dependencies, we evaluate candidate models on datasets with temporal relationships spanning thousands of measurements. We compare model performance with a total of sixteen other state-of-the-art sequence modeling approaches, including the precursor models LinOSS-IM and LinOSS-IMEX. Experimental design and hyperparameter spreads are kept consistent across all models to ensure fair comparison.
Table 2: Test accuracies averaged over five different seeds on UEA time-series classification datasets. The highest score is indicated in bold and the second highest is underlined. The dataset names are abbreviations of the following UEA datasets: EigenWorms (Worms), SelfRegulationSCP1 (SCP1), SelfRegulationSCP2 (SCP2), EthanolConcentration (Ethanol), Heartbeat, and MotorImagery (Motor).
Worms SCP1 SCP2 Ethanol Heartbeat Motor Avg
Seq. length 17,984 896 1,152 1,751 405 3,000
# of Classes 5 2 2 4 2 2
NRDE 83.9 $\pm$ 7.3 80.9 $\pm$ 2.5 53.7 $\pm$ 6.9 25.3 $\pm$ 1.8 72.9 $\pm$ 4.8 47.0 $\pm$ 5.7 60.6
NCDE 75.0 $\pm$ 3.9 79.8 $\pm$ 5.6 53.0 $\pm$ 2.8 29.9 $\pm$ 6.5 73.9 $\pm$ 2.6 49.5 $\pm$ 2.8 60.2
Log-NCDE 85.6 $\pm$ 5.1 83.1 $\pm$ 2.8 53.7 $\pm$ 4.1 34.4 $\pm$ 6.4 75.2 $\pm$ 4.6 53.7 $\pm$ 5.3 64.3
LRU 87.8 $\pm$ 2.8 82.6 $\pm$ 3.4 51.2 $\pm$ 3.6 21.5 $\pm$ 2.1 78.4 $\pm$ 6.7 48.4 $\pm$ 5.0 61.7
S5 81.1 $\pm$ 3.7 89.9 $\pm$ 4.6 50.5 $\pm$ 2.6 24.1 $\pm$ 4.3 77.7 $\pm$ 5.5 47.7 $\pm$ 5.5 61.8
S6 85.0 $\pm$ 16.1 82.8 $\pm$ 2.7 49.9 $\pm$ 9.4 26.4 $\pm$ 6.4 76.5 $\pm$ 8.3 51.3 $\pm$ 4.7 62.0
Mamba 70.9 $\pm$ 15.8 80.7 $\pm$ 1.4 48.2 $\pm$ 3.9 27.9 $\pm$ 4.5 76.2 $\pm$ 3.8 47.7 $\pm$ 4.5 58.6
LinOSS-IMEX 80.0 $\pm$ 2.7 87.5 $\pm$ 4.0 58.9 $\pm$ 8.1 29.9 $\pm$ 1.0 75.5 $\pm$ 4.3 57.9 $\pm$ 5.3 65.0
LinOSS-IM 95.0 $\pm$ 4.4 87.8 $\pm$ 2.6 58.2 $\pm$ 6.9 29.9 $\pm$ 0.6 75.8 $\pm$ 3.7 60.0 $\pm$ 7.5 67.8
D-LinOSS 93.9 $\pm$ 3.2 88.9 $\pm$ 3.0 58.6 $\pm$ 2.3 29.9 $\pm$ 0.6 75.8 $\pm$ 4.9 61.1 $\pm$ 2.0 68.0
4.1 UEA time-series classification
We consider a benchmark from the University of East Anglia (UEA) Multivariate Time Series Classification Archive (UEA-MTSCA) Bagnall et al. (2018) introduced in Walker et al. (2024). This benchmark consists of six datasets chosen to evaluate the ability of sequence models to capture long-range interactions. The UEA datasets are classification tasks, ranging from classifying organisms from motion readings (EigenWorms) to classifying fluid alcohol percentage based on measurements of transmissive light spectra (EthanolConcentration). We precisely follow the experimental design proposed in Walker et al. (2024), conducting a model hyperparameter search over a grid of 162 predetermined configurations for each dataset. Further, each model instance is trained on five seeds, and the average test accuracy for the top performing model instances are reported. The high scoring hyperparameter configurations of D-LinOSS model instances are tabulated in Appendix B.3. All models use the same 70-15-15 train-validation-test data splits, controlled by the seed for a given trial. Model scores for LinOSS-IM and LinOSS-IMEX are sourced from Rusch & Rus (2025) and all other model scores are sourced from Walker et al. (2024).
Out of all models tested, D-LinOSS achieves the highest average test accuracy across the six UEA datasets–raising the previous high score from 67.8% to 68.0%. Notably, D-LinOSS improves state-of-the-art accuracy on MotorImagery by $1.1\%$ and scores in the top two for five out of the six datasets. D-LinOSS also outperforms the combination of both preceding models: the average score-wise maximum between LinOSS-IM and LinOSS-IMEX is $67.9\%$, still shy of D-LinOSS. D-LinOSS improves on or matches the second best model, LinOSS-IM, in all but one dataset, EigenWorms, which is the smallest dataset out of the six.
4.2 PPG-DaLiA time-series regression
We evaluate model performance on the PPG dataset for motion compensation and heart rate estimation in Daily Life Activities (PPG-DaLiA) (Reiss et al., 2019), which is a time-series regression task. Here, models are challenged with learning human heart rate patterns as a function of various sensor measurements, such as ECG readings, wearable accelerometers, and respiration sensing. The dataset consists of 15 different subjects performing a variety of daily tasks, and bio-sensory data is collected in sequences up to 50,000 points in length. We follow the same experimental design as before, searching model hyperparameters over a grid of 162 configurations and training each model instance on five seeds. All models use the same 70-15-15 data split. D-LinOSS achieves the best results, reducing the lowest MSE from 6.4 to 6.16 ($\times 10^{-2}$) compared to LinOSS-IM.
Table 3: Test accuracies averaged over five different seeds on the PPG-DaLiA time-series regression dataset. The best score is indicated in bold and the second best is underlined.
Model MSE $\times 10^{-2}$
NRDE Morrill et al. (2021) 9.90 $\pm$ 0.97
NCDE (Kidger et al., 2020) 13.54 $\pm$ 0.69
Log-NCDE Walker et al. (2024) 9.56 $\pm$ 0.59
LRU Orvieto et al. (2023) 12.17 $\pm$ 0.49
S5 Smith et al. (2023) 12.63 $\pm$ 1.25
S6 Gu & Dao (2023) 12.88 $\pm$ 2.05
Mamba Gu & Dao (2023) 10.65 $\pm$ 2.20
LinOSS-IMEX Rusch & Rus (2025) 7.5 $\pm$ 0.46
LinOSS-IM Rusch & Rus (2025) 6.4 $\pm$ 0.23
D-LinOSS 6.16 $\pm$ 0.73
4.3 Weather time-series forecasting
To assess the generality of D-LinOSS as a sequence-to-sequence model, as in Gu et al. (2021), we evaluate its performance on a long-horizon time-series forecasting task without any architectural modifications. In this setting, forecasting is framed as a masked sequence transformation problem, allowing the model to predict future values based solely on partially masked input sequences.
We focus on the weather forecasting task introduced in Zhou et al. (2021), which involves predicting one month of hourly measured multivariate local climatological data based on the previous month’s measurements. The dataset spans 1,600 U.S. locations between 2010 to 2013; further details are provided in NCEI (2025).
In this benchmark, D-LinOSS is compared against Transformer-based architectures, LSTM variants, the structured state-space model S4, and previous versions of LinOSS. D-LinOSS achieves the best performance, reducing the lowest mean absolute error (MAE) from 0.508 (LinOSS-IMEX) to 0.486.
Table 4: Mean absolute error on the weather dataset predicting the future 720 time steps based on the past 720 time steps. The best score is indicated in bold and the second best is underlined.
Model Mean Absolute Error
Informer (Zhou et al., 2021) 0.731
Informer (Zhou et al., 2021) 0.741
LogTrans (Li et al., 2019) 0.773
Reformer (Kitaev et al., 2020) 1.575
LSTMa (Bahdanau et al., 2016) 1.109
LSTnet (Lai et al., 2018) 0.757
S4 (Gu et al., 2021) 0.578
LinOSS-IMEX Rusch & Rus (2025) 0.508
LinOSS-IM Rusch & Rus (2025) 0.528
D-LinOSS 0.486
4.4 Synthetic Adding
An additional practical benefit of D-LinOSS is faster training convergence compared to preceding models. To showcase this, we consider the synthetic adding task (Hochreiter, 1997), where the model must compute the sum of two randomly specified numbers embedded in a sequence of white noise. To make the task challenging, models are required to predict the sum using only the final output token, rather than aggregating all outputs across the sequence. We evaluate sequence lengths of 500, 2000, and 5000, and perform a small grid search for all LinOSS variants. Further details are provided in Appendix B.2.
[…]
Figure 2: Validation metric convergence for the adding task of different sequence lengths.
5 Related Work
State Space Models (SSMs) were introduced as a deep learning framework for sequential data in Gu et al. (2021). Early variants (Gu et al., 2022; Nguyen et al., 2022; Goel et al., 2022) relied on Fast Fourier Transform (FFT) and HiPPO parameterizations (Gu et al., 2020) to efficiently compute linear recurrences. More recent architectures employ diagonal state matrices in combination with fast associative parallel scans, which was first developed for RNNs (Martin & Cundy, 2017; Kaul, 2020) and later adapted to SSMs in Smith et al. (2023). While these models initially required HiPPO matrices to initialize weights, subsequent work has shown that simple random initialization is sufficient (Orvieto et al., 2023). Finally, D-LinOSS and the models above are based on LTI systems, there is growing interest in time-varying SSMs for challenging domains such as language and video (Gu & Dao, 2023; Dao & Gu, 2024; Hasani et al., 2022; Merrill et al., 2024).
The most closely related model to our proposed D-LinOSS is the original LinOSS, introduced in Rusch & Rus (2025). While LinOSS was the first SSM to explicitly leverage oscillatory dynamics, the idea of incorporating oscillatory behavior into deep learning architectures has also appeared in other domains. For instance, recurrent models such as coupled oscillatory RNNs (coRNNs) (Rusch & Mishra, 2021a) and UnICORNNs (Rusch & Mishra, 2021b) introduce oscillations into their hidden state dynamics, while graph-based approaches like Graph Coupled Oscillator Networks (GraphCON) (Rusch et al., 2022) extend similar principles to structured data.
6 Discussion and conclusion
We introduced D-LinOSS, an extension of the LinOSS model that incorporates learnable damping across arbitrary time scales. Through spectral analysis, we showed that existing LinOSS variants are rigidly defined by their discretization scheme and can only express a narrow set of dynamical behaviors. In contrast, D-LinOSS captures the full range of stable second-order dynamics.
[…]
While D-LinOSS demonstrates strong empirical results as a general sequence model, it is based on layers of LTI systems, which are fundamentally limited in their ability to capture certain contextual dependencies, such as selective copying (Gu & Dao (2023), Jing et al. (2019)). Building on the growing interest in time-varying SSMs sparked by Gu & Dao (2023); Dao & Gu (2024), we aim to explore future work on variants of D-LinOSS that integrate the selectivity of time-varying dynamics.
As D-LinOSS is inherently well-suited to represent temporal relationships with oscillatory structure, we aim to explore applications where such patterns are fundamental. In particular, climate science, seismic monitoring, and astrophysics data all exhibit complex patterns governed by oscillatory behavior. Moving forward, we believe that D-LinOSS will be a key player in advancing machine-learning based approaches in domains grounded in the physical sciences.
Acknowledgments
This work was supported in part by the Postdoc.Mobility grant P500PT-217915 from the Swiss National Science Foundation, the Schmidt AI2050 program (grant G-22-63172), and the Department of the Air Force Artificial Intelligence Accelerator and was accomplished under Cooperative Agreement Number FA8750-19-2-1000. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of the Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
The authors acknowledge the MIT SuperCloud (Reuther et al. (2018)) and Lincoln Laboratory Supercomputing Center for providing HPC resources that have contributed to the research results reported within this paper.
Supplementary Material for:
Learning to Dissipate Energy in Oscillatory State-Space Models
Appendix A Theoretical properties
A.1 D-LinOSS eigenvalues
Proposition A.1.
The eigenvalues of the D-LinOSS recurrent matrix $\mathbf{M}\in\mathbb{R}^{2m\times 2m}$ are
$\lambda_{i_{1,2}}=\frac{1+\frac{\Delta t_{i}}{2}\mathbf{G}_{i}-\frac{\Delta t_{i}^{2}}{2}\mathbf{A}_{i}}{1+\Delta t_{i}\mathbf{G}_{i}}\pm\frac{\frac{\Delta t_{i}}{2}\sqrt{(\mathbf{G}_{i}-\Delta t_{i}\mathbf{A}_{i})^{2}-4\mathbf{A}_{i}}}{1+\Delta t_{i}\mathbf{G}_{i}},$
where pairs of eigenvalues are denoted as $\lambda_{i_{1,2}}$ and i=1,2,,mi=1,2,...,m.
Proof.
$\mathbf{M}$ (6) is a matrix with diagonal sub-blocks $\mathbf{M}_{11}$, $\mathbf{M}_{12}$, $\mathbf{M}_{21}$, and $\mathbf{M}_{22}$, i.e. it follows the structure:
$\mathbf{M}=\begin{vmatrix}\begin{bmatrix}(\mathbf{M}_{11})_{1}&&\\ &\ddots&\\ &&(\mathbf{M}_{11})_{m}\end{bmatrix}&\begin{bmatrix}(\mathbf{M}_{12})_{1}&&\\ &\ddots&\\ &&(\mathbf{M}_{12})_{m}\end{bmatrix}\\ \\ \begin{bmatrix}(\mathbf{M}_{21})_{1}&&\\ &\ddots&\\ &&(\mathbf{M}_{21})_{m}\end{bmatrix}&\begin{bmatrix}(\mathbf{M}_{22})_{1}&&\\ &\ddots&\\ &&(\mathbf{M}_{22})_{m}\end{bmatrix}\end{vmatrix}$
Since $\mathbf{M}$ represents $m$ uncoupled oscillatory systems, we see that the operation $\mathbf{M}\mathbf{w}$, $\mathbf{w}^{\intercal}=[\mathbf{z}^{\intercal},\mathbf{x}^{\intercal}]$ can be independently split across the $m$ pairs of state variables $\mathbf{z}_{i},\mathbf{x}_{i}]$. Thus, it suffices to re-arrange $\mathbf{M}$ into $m$ different $2\times 2$ systems and analyze the eigenvalues of each system separately. Substituting the real expressions for $\mathbf{M}_{11}$, $\mathbf{M}_{12}$, $\mathbf{M}_{21}$, and $\mathbf{M}_{22}$, and using brackets to denote the $2\times 2$ matrix formed by diagonally indexing along the sub-blocks of $\mathbf{M}$, the $i^{th}$ system matrix is:
$\mathbf{M}]_{i}=\begin{bmatrix}\mathbf{S}^{-1}_{i}&-\Delta t_{i}\mathbf{S}^{-1}_{i}\mathbf{A}_{i}\\ \Delta t_{i}\mathbf{S}^{-1}_{i}&1-\Delta t_{i}^{2}\mathbf{S}^{-1}_{i}\mathbf{A}_{i}\end{bmatrix}$
A straightforward exercise using the definition $\mathbf{S}=\mathbf{I}+\Delta t\mathbf{G}$ leads us to the $i^{th}$ pair of eigenvalues.
$\det(\lambda\mathbf{I}-[\mathbf{M}]_{i})$ $=\det\begin{bmatrix}\lambda-\mathbf{S}^{-1}_{i}&\Delta t_{i}\mathbf{S}^{-1}_{i}\mathbf{A}_{i}\\ -\Delta t_{i}\mathbf{S}^{-1}_{i}&\lambda-1+\Delta t_{i}^{2}\mathbf{S}^{-1}_{i}\mathbf{A}_{i}\end{bmatrix}$
$=(\lambda-\mathbf{S}^{-1}_{i})(\lambda-1+\Delta t_{i}^{2}\mathbf{S}^{-1}_{i}\mathbf{A}_{i})+(\Delta t_{i}\mathbf{S}^{-1}_{i})(\Delta t_{i}\mathbf{S}^{-1}_{i}\mathbf{A}_{i})$
$=\lambda^{2}+\lambda(\mathbf{S}^{-1}_{i}(\Delta t^{2}_{i}\mathbf{A}_{i}-1)-1)+\mathbf{S}^{-1}_{i}$
$=\lambda^{2}+\lambda\frac{-2-\Delta t_{i}\mathbf{G}_{i}+\Delta t_{i}^{2}\mathbf{A}_{i}}{1+\Delta t_{i}\mathbf{G}_{i}}+\frac{1}{1+\Delta t_{i}\mathbf{G}_{i}}=0$
$\implies\lambda_{i_{1,2}}$ $=\frac{1+\frac{\Delta t_{i}}{2}\mathbf{G}_{i}-\frac{\Delta t_{i}^{2}}{2}\mathbf{A}_{i}}{1+\Delta t_{i}\mathbf{G}_{i}}\pm\frac{\frac{\Delta t_{i}}{2}\sqrt{(\mathbf{G}_{i}-\Delta t_{i}\mathbf{A}_{i})^{2}-4\mathbf{A}_{i}}}{1+\Delta t_{i}\mathbf{G}_{i}}$
A.2 Stability criterion
Proposition A.2.
Assume $\mathbf{A}_{i}$, $\mathbf{G}_{i}$ are non-negative, and $\Delta t_{i}\in(0,1]$. If the following is satisfied:
$(\mathbf{G}_{i}-\Delta t_{i}\mathbf{A}_{i})^{2}\leq 4\mathbf{A}_{i}$ (11)
Then $\lambda_{i_{1,2}}$ come in complex conjugate pairs $\lambda_{i}$, $\lambda_{i}^{*}$ with the following magnitude:
$|\lambda_{i}|=\frac{1}{\sqrt{1+\Delta t_{i}\mathbf{G}_{i}}}\leq 1,$
i.e. the eigenvalues are unit-bounded. Define $\mathcal{S}_{i}$ to be the set of all $(\mathbf{A}_{i},\mathbf{G}_{i})$ that satisfy the above condition. For notational convenience, we order the eigenvalues such that $\text{\emph{Im}}(\lambda_{i})\geq 0$, $\text{\emph{Im}}(\lambda_{i}^{*})\leq 0$.
Proof.
$(\mathbf{G}_{i}-\Delta t_{i}\mathbf{A}_{i})^{2}-4\mathbf{A}_{i}$ is the determinant of each eigenvalue pair, so $(\mathbf{G}_{i}-\Delta t_{i}\mathbf{A}_{i})^{2}\leq 4\mathbf{A}_{i}$ means $\lambda_{i}$ can be written in complex form with the following real and imaginary components.
$\text{Re}(\lambda_{i})$ $=\frac{1+\frac{\Delta t_{i}}{2}\mathbf{G}_{i}-\frac{\Delta t_{i}^{2}}{2}\mathbf{A}_{i}}{1+\Delta t_{i}\mathbf{G}_{i}}$
$\text{Im}(\lambda_{i})$ $=\pm\frac{\frac{\Delta t_{i}}{2}\sqrt{4\mathbf{A}_{i}-(\mathbf{G}_{i}-\Delta t_{i}\mathbf{A}_{i})^{2}}}{1+\Delta t_{i}\mathbf{G}_{i}}$
The magnitude of this complex number is:
$|\lambda_{i}|=\sqrt{\text{Re}(\lambda_{i})^{2}+\text{Im}(\lambda_{i})^{2}}=\frac{1}{\sqrt{1+\Delta t_{i}\mathbf{G}_{i}}}$
$\Delta t_{i},\mathbf{G}_{i}\geq 0\implies|\lambda_{i}|\leq 1$
We note that this stability criterion is a sufficient but not necessary condition. There exists solutions $(\mathbf{A}_{i},\mathbf{G}_{i})$ rendering $|\lambda_{i}|\leq 1$ that do not lie in $\mathcal{S}_{i}$. However, as shown in Proposition 3.3, there already exists a bijection from $\mathcal{S}_{i}$ to the full complex unit disk.
A.3 Spectral image of D-LinOSS
Proposition A.3.
The mapping $\Phi:\mathcal{S}_{i}\rightarrow\mathbb{C}_{|z|\leq 1}$ defined by $(\mathbf{A}_{i},\mathbf{G}_{i})\mapsto(\lambda_{i},\lambda_{i}^{*})$ is bijective.
Proof.
Assuming constant $\Delta t_{i}\in(0,1]$, it can be checked via symbolic manipulation that the following expression is the inverse of the eigenvalues from Proposition 3.1:
$\mathbf{A}_{i}=\frac{\lambda_{i}\lambda^{*}_{i}-\lambda_{i}-\lambda^{*}_{i}+1}{\Delta t_{i}^{2}\lambda_{i}\lambda^{*}_{i}},\ \mathbf{G}_{i}=\frac{1-\lambda_{i}\lambda^{*}_{i}}{\Delta t_{i}\lambda_{i}\lambda^{*}_{i}}$
Proposition 3.2 has already shown that $\Phi(\mathcal{S}_{i})\subseteq\mathbb{C}_{|z|\leq 1}$, so it remains to show that $\Phi^{-1}(\mathbb{C}_{|z|\leq 1})\subseteq\mathcal{S}_{i}$ to conclude that the mapping is bijective between $\mathcal{S}_{i}$ and $\mathbb{C}_{|z|\leq 1}$.
It’s clear that $\mathbf{A}_{i}$, $\mathbf{G}_{i}$ are non-negative as the numerators are denominators of the above expressions are both non-negative given $|\lambda_{i}|\leq 1$. The final condition simplifies to:
$(\mathbf{G}_{i}-\Delta t_{i}\mathbf{A}_{i})^{2}-4\mathbf{A}_{i}\leq 0\iff\frac{(\lambda_{i}-\lambda_{i}^{*})^{2}}{(\Delta t_{i}\lambda_{i}\lambda_{i}^{*})^{2}}=\frac{-4\text{Im}(\lambda_{i})^{2}}{(\Delta t_{i}\lambda_{i}\lambda_{i}^{*})^{2}}\leq 0$
which is also true.
Figure 3: Reachable eigenvalue sets for the different oscillatory SSMs. The eigenvalue $\lambda=0.8$ of the exponential decay experiment from Section 3.2 only lies in the spectral range of D-LinOSS.
A.4 Set measure of LinOSS eigenvalues
Proposition A.4.
For both LinOSS-IM and LinOSS-IMEX, the set of eigenvalues constructed from $\mathbf{A}_{i}\in\mathbb{R}_{\geq 0}$ and $\Delta t_{i}\in(0,1]$ is of measure zero in $\mathbb{C}$.
Proof.
Recall the eigenvalue expressions from Rusch & Rus (2025):
$\lambda^{\text{IM}}_{i_{1,2}}=\frac{1}{1+\Delta t_{i}^{2}\mathbf{A}_{i}}\pm j\frac{\Delta t_{i}\sqrt{\mathbf{A}_{i}}}{1+\Delta t_{i}^{2}\mathbf{A}_{i}},\quad\lambda^{\text{IMEX}}_{i_{1,2}}=\frac{1}{2}(2-\Delta t_{i}^{2}\mathbf{A}_{i})\pm\frac{j}{2}\sqrt{\Delta t_{i}^{2}\mathbf{A}_{i}(4-\Delta t_{i}^{2}\mathbf{A}_{i})},$
Since $\lambda_{i_{1}}=\lambda_{i_{2}}^{*}$, it suffices to prove the proposition for the first eigenvalue, i.e. $\{\lambda_{i_{1}}\in\mathbb{C}\ |\ \mathbf{A}_{i}\in\mathbb{R}_{\geq 0},\Delta t_{i}\in(0,1]\}$ is a set of measure zero. We start with the following lemma:
Lemma A.5.
Let $f:M\rightarrow N$ be a continuously differentiable map of manifolds where $dimM<dimN$. Then $f(M)$ is of measure zero in N (Gualtieri, 2016).
Using Lemma A.5, it suffices to show that range($\lambda_{i_{1}}^{\text{IM}}$) is a 1-manifold in $\mathbb{C}\cong\mathbb{R}^{2}$. Beginning with the LinOSS-IM expression, apply the change of variables $\gamma_{i}:=\Delta t_{i}\sqrt{\mathbf{A}_{i}}$, where $\mathbf{A}_{i}\in\mathbb{R}_{\geq 0},\Delta t_{i}\in(0,1]\iff\gamma_{i}\in\mathbb{R}_{\geq 0}$:
$\lambda_{i_{1}}^{\text{IM}}(\gamma_{i})=\bigg(\frac{1}{1+\gamma_{i}^{2}},\frac{\gamma_{i}}{1+\gamma_{i}^{2}}\bigg)$
This map is continuously differentiable, injective on $\mathbb{C}$ (and surjective onto its image), and its inverse is continuously differentiable, so it satisfies the conditions of Lemma A.5. So, the range of $\lambda_{i_{1}}^{\text{IM}}$ is an embedded 1-manifold in $\mathbb{R}^{2}$ and therefore a set of measure zero.
Using the same change of variables, an identical argument can be applied to the LinOSS-IMEX expression to show it also produces a set of measure zero; this argument is omitted for concision.
A.5 Universality
An operator $\Phi:C_{0}([0,T];\mathbb{R}^{p})\rightarrow C_{0}([0,T];\mathbb{R}^{q})$ is said to be causal if $\forall t\in[0,T]$, if $\mathbf{u},\mathbf{v}\in C_{0}([0,T];\mathbb{R}^{p})$ are two input signals such that $\mathbf{u}|_{[0,t]}\equiv\mathbf{v}|_{[0,t]}$, then $\Phi(\mathbf{u})(t)=\Phi(\mathbf{v})(t)$.
An operator is said to be continuous if $\Phi:(C_{0}([0,T];\mathbb{R}^{p}),\|{\,\cdot\,}\|_{\infty})\to(C_{0}([0,T];\mathbb{R}^{q}),\|{\,\cdot\,}\|_{\infty})$, i.e. $\Phi$ is a map between continuous functions with respect to the $L^{\infty}$-norms on the input/output signals.
The theorem of Rusch & Rus (2025) is briefly paraphrased:
Theorem A.6.
Let $\Phi$ be any causal and continuous operator. Let $K\subset C_{0}([0,T];\mathbb{R}^{p})$ be compact. Then for any $\epsilon>0$, there exists a configuration of hyperparameters and weight matrices, such that the output $\mathbf{z}:[0,T]\to\mathbb{R}^{q}$ of a LinOSS block satisfies:
$\sup_{t\in[0,T]}|\Phi(\mathbf{u})(t)-\mathbf{z}(t)|\leq\epsilon,\quad\forall\,\mathbf{u}\in K.$
In other words, a LinOSS block can approximate any causal and continuous operator on compact input spaces to arbitrarily high accuracy.
Appendix B Experiments and results
B.1 Regression experiment
A small hyperparameter grid search was conducted over hidden dimension $\in\{8,64\}$, SSM dimension $\in\{8,64\}$, and number of blocks $\in\{2,6\}$. Learning rate was kept constant at 1e-3.
Input sequences were sampled from white noise and passed through the discrete state-space system corresponding to parameters $A=0.8,B=1,C=1,D=0$. A sequence length of $1000$ was selected to study the model behaviors within the regime of long range learning.
B.2 Adding task
Models were varied across number of blocks $\in\{2,4,6\}$ while keeping constant hidden dimension =128=128, SSM dimension =128=128, and learning rate == 1e-4. Each model configuration was trained on 5 different seeds and the highest performing model instances (as measured by fastest validation set convergence time) are compared in Figure 2.
B.3 Hyperparameters
[…]
Table 5: Best performing hyperparameters for D-LinOSS across each of the eight datasets.
Dataset lr hidden dim state dim num blocks include time
Worms 1e-3 128 64 2 False
SCP1 1e-4 128 256 6 True
SCP2 1e-5 128 64 6 False
Ethanol 1e-5 16 256 4 False
Heartbeat 1e-4 16 16 2 False
Motor 1e-3 16 64 4 False
PPG 1e-3 64 64 4 True
Weather 7.95e-5 128 128 3 False
B.4 Compute requirements
Below, we tabulate the compute resources required for each model across all datasets considered in the UEA-MTSCA classification experiments of Section 4.1. The main table is sourced from Rusch & Rus (2025), which lists the total number of parameters, average GPU memory usage measured in MB, and average run time per 1000 training steps measured in seconds. All models operate on the same codebase and python libraries, adopted from both Walker et al. (2024) and Rusch & Rus (2025). These compute requirements are evaluated using an Nvidia RTX 4090 GPU.
Table 6: Compute requirements for the classification experiments considered in Section 4.1.
NRDE NCDE Log-NCDE LRU S5 Mamba S6 LinOSS-IMEX LinOSS-IM D-LinOSS
Worms $\lvert\theta\rvert$ 105110 166789 37977 101129 22007 27381 15045 26119 134279 134279
mem. 2506 2484 2510 10716 6646 13486 7922 6556 3488 3488
time 5386 24595 1956 94 31 122 68 37 14 14
SCP1 $\lvert\theta\rvert$ 117187 166274 91557 25892 226328 184194 24898 447944 991240 992776
mem. 716 694 724 960 1798 1110 904 4768 4772 4790
time 1014 973 635 9 17 7 3 42 38 38
SCP2 $\lvert\theta\rvert$ 200707 182914 36379 26020 5652 356290 26018 448072 399112 399496
mem. 712 692 714 954 762 2460 1222 4772 2724 2736
time 1404 1251 583 9 9 32 7 55 22 22
Ethanol $\lvert\theta\rvert$ 93212 133252 31452 76522 76214 1032772 5780 70088 6728 71112
mem. 712 692 710 1988 1520 4876 938 4766 1182 4774
time 2256 2217 2056 16 9 255 4 48 8 37
Heartbeat $\lvert\theta\rvert$ 15657742 1098114 168320 338820 158310 1034242 6674 29444 10936 4356
mem. 6860 1462 2774 1466 1548 1650 606 922 928 672
time 9539 1177 826 8 11 34 4 4 7 4
Motor $\lvert\theta\rvert$ 1134395 186962 81391 107544 17496 228226 52802 106024 91844 20598
mem. 4552 4534 4566 8646 4616 3120 4056 12708 4510 4518
time 7616 3778 730 51 16 35 34 128 11 20
B.5 Initialization techniques
Figure 4: Initialization study varying intervals of eigenvalue magnitude and methods of sampling.
To understand how to best initialize the D-LinOSS parameters, we conduct a study on model performance comparing four different sampling techniques.
  • Experiment 1: Uniformly sample $\mathbf{G}\in[0,G_{\max}]$ for different values of $G_{\max}$ and uniformly sample $\mathbf{A}\in[0,1]$
  • Experiment 2: Uniformly sample $\mathbf{G}\in[0,G_{\max}]$ for different values of $G_{\max}$ and uniformly sample $\phi\in[0,\pi)$
  • Experiment 3: Radially uniform sample $|\lambda|\in[r_{\min},1]$ for different values of $r_{\min}$ and uniformly sample $\mathbf{A}\in[0,1]$
  • Experiment 4: Radially uniform sample $|\lambda|\in[r_{\min},1]$ for different values of $r_{\min}$ and uniformly sample $\phi\in[0,\pi)$
Where sampling $\mathbf{G}$ or sampling $|\lambda|$ are two different ways of controlling eigenvalue magnitude, and sampling $\mathbf{A}$ or sampling $\phi:=\arg(\lambda)$ are two different ways of controlling eigenvalue phase. Note that eigenvalues come in complex conjugate pairs, so sampling a single eigenvalue in $\phi\in[0,\pi)$ covers the full complex disk. “Radially uniform sampling" refers to uniformly sampling over the area of the cartesian coordinate ring described by the polar coordinate boundaries as opposed to uniformly sampling in the polar coordinates themselves.
For each experiment, values of $G_{\max}$ or $r_{\min}$ are varied to generate different eigenvalue distributions. For each evaluation of an initialization technique (a single point on figure 4), D-LinOSS is trained over a small sweep of 16 hyper-parameter configurations, and each model configuration is trained on 5 different seeds. This process is repeated for the three datasets SelfRegulationSCP1, Heartbeat, and MotorImagery. Figure 4 displays the average highest test score over all three datasets.
Radially uniform sampling eigenvalue magnitude and uniformly sampling eigenvalue phase is the best performing initialization technique observed. For these three datasets, a lower bound of $r_{\min}=0.9$ results in the highest average test accuracy.